OpenStudy (anonymous):

A Wire of length 10 meter is cut into 2 portions ,One of them is bent to form a circle and the other is bent to form a square . Find the radius of the circle and the side of the square when the sum of the areas of the circles and the square is minimum .

OpenStudy (anonymous):

@satellite73 : May u check this ?

OpenStudy (anonymous):

Do you have any ideas?

OpenStudy (anonymous):

well,My attempt was : 1-I got the Area. 2-I differentiate dA/dx with respect to dA/dx =0 and then i stopped xD

OpenStudy (anonymous):

alright what did you get the area function

OpenStudy (anonymous):

Your restriction should be \[x+y=10\]

OpenStudy (anonymous):

Its not about the area function ,Its about the idea ..I dont even know if i started it right or not :/

OpenStudy (anonymous):

but that's perimeter and you're looking for area so you need to relate the two

OpenStudy (anonymous):

well you're minimizing the area so you should be minimizing the area of a circle and an area of a rectangle

OpenStudy (anonymous):

your restrictionis that you only have 10 meters of wire

OpenStudy (anonymous):

That is the function you need ^ he related the perimeter and circumference to area. Now if you find the min area, you can solve for r

OpenStudy (anonymous):

wait mertsj, isn't 10-2pir what is left after the circle meaning the parameter of the square is 10-2pir

OpenStudy (anonymous):

perimeter*

OpenStudy (anonymous):

A=A1+A2 \[\Large A=\frac{1}{16}X^{2} + \frac{1}{4\pi} (10-x)^{2}\]

OpenStudy (anonymous):

if you have the radius as \(r\) then the area of the circle \(\pi r^2\) and the circumference is \(2\pi r\) so each side of the square is \[\frac{10-2\pi r}{4}\] and its area is \[\left(\frac{10-2\pi r}{4}\right)^2\]

OpenStudy (anonymous):

total area is therefore \[\pi r^2+\left(\frac{10-2\pi r}{4}\right)^2\] and that is what you are trying to maximize

OpenStudy (anonymous):

minimize

OpenStudy (anonymous):

He need the radius and the side ,When the circle and square are minimized .

OpenStudy (anonymous):

OMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMG , I GOT IT !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! TY GUYS TY SOO MUCH !

OpenStudy (anonymous):

The secret on the word 2 portions ,and When the circle and square are minimized .

OpenStudy (anonymous):

oh minimized, sorry but the same idea right? and when you solve for \(r\) you still need to solve for the length that you use for the square, which should be easy enough since it is \(10-2\pi r\)

OpenStudy (anonymous):

sattellite please help me

OpenStudy (anonymous):

its about second derivative though ..

OpenStudy (anonymous):

i cheated minimum is at \(r=\frac{4}{5+\pi}\) http://www.wolframalpha.com/input/?i= \pi+r^2%2B\left%28\frac{10-2\pi+r}{4}\right%29^2

OpenStudy (anonymous):

the answer is 5/Pi+4 . Ty for your help :)